OpenStudy (shamil98):

derivative of x^2cosx product rule?

hartnn (hartnn):

yep

OpenStudy (shamil98):

f'(x)g(x) + f(x)g'(x) f(x) = x^2 f'(x) = 2x g(x) = cosx g'(x) = -sinx so.. 2x(cosx) + x^2(-sinx) ?

hartnn (hartnn):

correct!

OpenStudy (shamil98):

yay :D